test the series : infinity sigma n=1 (-1) ^n-1 sin nx/n√n for absolute and conditional convergence

1
Expert's answer
2022-02-15T15:14:45-0500

"\Sigma_{n=1}^{\infty} \dfrac{(-1)^{n-1}\sin nx}{n\sqrt n}
\\=\Sigma_{n=1}^{\infty} \dfrac{(-1)^{n-1}\sin nx}{n^{3/2}}"

Let "a_n=\dfrac{(-1)^{n-1}\sin nx}{n^{3/2}}"

"\Rightarrow a_{n+1}=\dfrac{(-1)^{n}\sin (n+1)x}{(n+1)^{3/2}}"

Now, "p=\lim_{n\rightarrow \infty}|\dfrac{a_{n+1}}{a_n}|=\lim_{n\rightarrow \infty}|\dfrac{\dfrac{(-1)^{n}\sin (n+1)x}{(n+1)^{3/2}}}{\dfrac{(-1)^{n-1}\sin nx}{n^{3/2}}}|"

"=\lim_{n\rightarrow \infty}|{\dfrac{(-1)\sin (n+1)x.n^{3/2}}{(n+1)^{3/2}\sin nx}}|
\\=\lim_{n\rightarrow \infty}|{\dfrac{\sin (n+1)x}{(1+\frac 1n)^{3/2}\sin nx}}|
\\=\lim_{n\rightarrow \infty}|{\dfrac{1}{(1+\frac 1n)^{3/2}}}|\times \lim_{n\rightarrow \infty}|{\dfrac{\sin (n+1)x}{\sin nx}}|
\\=1\times \infty
\\= \infty"

So, "p>1"

Thus, the series is divergent by the ratio test.